OpenStudy (anonymous):

what is the equation to the line tangent to h(x)=2x^2-4x+5 through the point (1,3)... I know the derivative is the slope and that is 4x-4.. idk what to do next though :\

OpenStudy (radar):

Then the slope would be 4x-4 or 4(1)-4 or 0 0=(y-3)/x-1 y-3=0 y=3 Looks like a horizontal line y=3